A stone is tossed straight upwards from a height of 2 meters with an initial positive velocity


of 30 m/s. If time it takes to hit the ground is 10 seconds, what is the maximum height it


reaches from the ground?

1
Expert's answer
2022-06-06T08:53:30-0400
"h=\frac{v_0^2}{2g}=\frac{30^2}{2*10}=45\:\rm m"

"h_{\max}=h+h_0=45+2=47\:\rm m"